1.00 x 10^3 cL = ________L
1
Expert's answer
2020-10-02T14:19:42-0400

"cL" represents centi litres.

"1L=100cL\implies 1\ cL=10^{-2}L" .

Using the above relation,

"1.00\times 10^3\ cL =1.00\times 10^3 \times 10^{-2}\ L=1.00\times 10^1" "L=10.00\ L"